OpenStudy (anonymous):

Using complete sentences, explain which method you would use to solve the following system of equations and why you chose that method. Show your work and provide the solution to the system. 3x + 2y = 7 4x – 3y = –2

OpenStudy (anonymous):

ohh i hate these questions like u can tell they r staright out of a textbook like idk they r soo annoying. ok fisrtly lets choose a method

OpenStudy (anonymous):

i would use elimination in my opinion

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so lets solve using elimination and eliminate they y so we wld have to multiply equation #1 by 3 and equation # 2 by 2

OpenStudy (anonymous):

3(3x+2y=7)= 9x+6y=21 2(4x-3y=-2)= 8x-6y= -4

OpenStudy (anonymous):

so we can add the 2 new equations and we get 17x=17 x=1

OpenStudy (anonymous):

mhm

OpenStudy (anonymous):

ok now we plug x into any of our equations 3(1)+2y=7 2y=4 y=2 x=1 and y=2 So these two equations meet at the point (1,2)

OpenStudy (anonymous):

I have chosen the elimination method since when using subsitution you are likely to end up with fractions which confuses me and I tend to make mistakes hahahhahahah

OpenStudy (anonymous):

idk i guess smth along those lines

OpenStudy (anonymous):

Well that is y i chose the elimmination method

OpenStudy (anonymous):

Since if i wld use the subsitution method I wld have to deal with fractions which I definately hate

OpenStudy (anonymous):

yeah me tooo! i feel yaa! thankss :)

OpenStudy (anonymous):

hah ur welcome